# LaM-SLidE: Latent Space Modeling of Spatial Dynamical Systems via Linked Entities

### v3

@@ 3.4 Training Procedure
The training process follows a two-stage approach similar to latent diffusion models [77]. First Stage: We train the encoder $\mathcal{E}$ and decoder $\mathcal{D}$, to reconstruct entities from latent space using assigned IDs, see Figure 2). Second Stage: We train the approximator $\mathcal{A}$ on the latent system representations produced by the frozen encoder $\mathcal{E}$ (details in Sections E.3 and 5).
@@ 4 Experiments
In order to evaluate LaM-SLidE we focus on three key aspects: (i) Robust generalization in diverse domains. We examine LaM-SLidE ’s generalization in different data domains in relation to other methods, for which we utilize tracking data from human motion behavior, trajectories of particle systems and data from molecular dynamics (MD) simulations. (ii) Temporal adaptability. We evaluate temporal adaptability through various conditioning/prediction horizons, considering single/multi-frame conditioning and short/long-term predictions; (iii) Computational efficiency and scalability. Finally, we assess inference time of LaM-SLidE , and evaluate performance with respect to model size. The subsequent sections show our key finding, for implementation details and additional results see App. E and G, for information on the datasets see Section E.1.
@@
For the MD experiments on peptides (Tetrapeptides), we use Jensen-Shannon divergence (JSD), evaluating the distribution of torsion angles, considering both, backbone (BB) and side chain (SC) angles. In order to capture long temporal behavior, we use Time-lagged Independent Component Analysis (TICA) [70], focusing on the slowest components TIC 0 and TIC 1. To investigate metastable state transitions we make use of Markov State Models (MSMs) [74, 65].
For inference time and scalability, we assess the number of function evaluations (NFE) and report performance of our method for different model sizes. A comprehensive overview of the conditioning and prediction horizons for each experiment is shown in Table 7.
@@ 4.1 Pedestrian Trajectory Forecasting (ETH-UCY)
For human motion behavior, we first consider the ETH-UCY dataset [68, 53], which provides pedestrian movement behavior, over five different scenes: ETH, Hotel, Univ, Zara1 and Zara2. We use the same setup as Han et al. [35], Xu et al. [93, 94], in which the methods obtain the first 8 frames as conditioning information and have to predict the next 12 frames. We report the minADE/ minFDE, computed across 20 sampled trajectories.
@@ Results
As shown in Table 2, our model performs competitively across all five scenes, achieving lower minFDE for Zara1 and Hotel scene, and lowest minADE on the ETH scene. Notably, in contrast to the compared baselines, we did not create additional features like velocity and acceleration or imply any kind of connectivity between entities.
@@ 4.2 Basketball Player Trajectory Forecasting (NBA)
Table 3: Results on the NBA dataset: Compared methods have to predict player positions for 12 frames and are given the initial 8 frames as input. Results in terms of minADE/minFDE for the Rebounding and Scoring scenarious.
@@
Our second experiment examines human motion behavior in the context of basketball gameplay. We use the SportVU NBA movement dataset [98], containing player movement data from the 2015-2016 NBA season. Each recorded frame includes ten player positions (5 for each team) and the ball position. We consider two different scenarios: a) Rebounding: containing scenes of missed shots; b) Scoring: containing scenes of a team scoring a basket. The interaction patterns – both in frequency and in their adversarial versus cooperative dynamics – exhibit different challenges as those in Section 4.1. The evaluation procedure by Xu et al. [94], which we use, provides 8 frames as input conditioning and the consecutive 12 frames for prediction. The performances is reported by the minADE/minFDE metrics, which are computed across 20 sampled trajectories. For the reported metrics, only the trajectories of the players are considered.
@@ Results
As illustrated in Table 3, our model shows robust performance across both scenarios, Rebounding and Scoring. For the Scoring scenario, LaM-SLidE achieves parity with SocialVAE [34] in terms of minADE and surpassing the performance in terms of minFDE. In the Rebounding scenario, we observe comparable but slightly lower performance of LaM-SLidE compared to SocialVAE. Trajectories sampled by our model are shown in Figure 9.
@@ 4.3 N-Body System Dynamics (Particle Systems)
We evaluate our method across three distinct N-Body simulation scenarios: a) Charged Particles: comprising particles with randomly assigned charges +1/1+1/-1+ 1 / - 1 interacting via Coulomb forces [51, 82]; b) Spring Dynamics: consisting of $N=5$ particles with randomized masses connected by springs with a probability $0.5$ between particle pairs [51]; and c) Gravitational Systems: containing $N=10$ particles with randomized masses and initial velocities governed by gravitational interactions [17]. For all three scenarios, we consider 10 conditioning frames and 20 frames for prediction. In line with Han et al. [35], we use $3000$ trajectories for training and $2000$ trajectories for validation and testing, we report ADE/FDE averaged over $K=5$ runs.
@@ Results
LaM-SLidE achieves the best performance in terms of ADE/FDE for the Charged Particles and Gravity scenarios and competitive second-rank performance in the Spring Dynamics scenario, see Table 2. Unlike compared methods, LaM-SLidE achieves these results without computing intermediate physical quantities such as velocities or accelerations. We present sampled trajectories in Figure 10.
@@ 4.4 Molecular Dynamics - Small Molecules (MD17)
In this experiment, we evaluate LaM-SLidE on the well established MD17 [21] dataset, containing simulated molecular dynamics trajectories of 8 small molecules. The size of those molecules ranges from 9 atoms (Ethanol and Malonaldehyde) to 21 atoms (Aspirin). We use 10 frames as conditioning, 20 frames for prediction and report ADE/FDE averaged over $K=5$ runs.
@@
Similar to Section 4.3, we compared LaM-SLidE to: Tensor Field Network [88], Radial Field [52], SE(3)-Transformer [30], EGNN [82], EqMotion [93], GeoTDM [35], and SVAE [94].
@@ Results
The results in Table 4 show the performance on the MD17 benchmark. LaM-SLidE achieves the lowest ADE/FDE of all methods and for all molecules. These results are particularly remarkable considering that: (1) our model operates without explicit definition molecular bond information, and (2) it surpasses the performance of all equivariant baselines, an inductive bias we intentionally omitted in LaM-SLidE .
@@
Table 4: Results on the MD17 dataset. Compared methods have to predict atom positions of 20 frames, conditioned on 10 input frames. Results in terms of ADE/FDE, averaged over 5 runs.
@@ 4.5 Molecular Dynamics - Tetrapeptides (4AA)
To investigate LaM-SLidE on long prediction horizons, we utilize the Tetrapeptide dataset from Jing et al. [44], containing explicit-solvent molecular dynamics trajectories simulated using OpenMM [26]. The dataset comprises 3,109 training, 100 validation and 100 test peptides. We use a single conditioning frame to predict 10,000 consecutive frames. The predictions are structured as a sequence of ten cascading 1,000-step rollouts, where each subsequent rollout is conditioned on the final frame of the previous. Note that, in contrast to the MD17 dataset, the methods predict trajectories of unseen molecules.
@@ Results
Table 5 shows performance metrics of the methods (for details on those metrics see Section E.6). Figure 12 shows the distribution of backbone torsions angles, and the free energy surfaces of the first two TICA components, for ground truth vs simulated trajectories. LaM-SLidE performs competitively with the current state-of-the-art method MDGen with respect to torsion angles, which is a notable achievement given that MDGen operates in torsion space only. With respect to the TICA and MSM metrics, LaM-SLidE even outperforms MDGen. Sampled trajectories are shown in Figure 13.
Table 5: Results on the Tetrapeptide dataset: Columns denote the JSD between distributions of torsion angles (backbone (BB), side-chain (SC), and all angles), the TICA, and the MSM metric.
@@ 4.6 Computational Efficiency and Scaling Behavior
To assess computational efficiency, we compare the NFEs of our model to the second-best method, GeoTDM. Our results show that LaM-SLidE requires up to 10x-100x fewer function evaluations depending on the domain. For a detailed discussion see Section G.2.
We further assess the scalability of LaM-SLidE across different model sizes. Our results indicate that performance consistently improves with increasing parameter count, suggesting that our method benefits from larger model capacity and could potentially achieve even better results with additional computational resources. Comprehensive details of this analysis can be found in App. G.
@@ 5 Discussion
Our experiments indicate that our architecture is applicable to a diverse set of problems; however, a few limitations provide opportunities for future improvement. While our current approach successfully allows to compress entities with beneficial reconstruction performance, our experiments indicate a tradeoff between the number of latent space vectors to encode system states and the performance of our latent model, for more details see Section G.3.
@@ Conclusion
We have introduced LaM-SLidE, a novel approach for modeling spatial dynamical systems consisting of a variable number of entities within a fixed-size latent system representation, where assignable identifiers allow traceability of individual entities. Across diverse domains, LaM-SLidE matches or exceeds specialized methods and offers promising scalability properties. It’s minimal reliance on prior knowledge makes it suitable for many tasks, suggesting its potential as a foundational architecture for dynamical systems.
@@ Acknowledgments
The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We thank NXAI GmbH, Audi AG, Silicon Austria Labs (SAL), Merck Healthcare KGaA, GLS (Univ. Waterloo), TÜV Holding GmbH, Software Competence Center Hagenberg GmbH, dSPACE GmbH, TRUMPF SE + Co. KG. We further thank Niklas Schmidinger, Anna Zimmel, Benedikt Alkin and Arturs Berzins for helpful discussions and feedback.
@@ Appendix A Notation
[…]
@@
[…]
@@
[…]
@@ B.1 Architecture Overview in Detail
Figure 5 shows an expanded view of our architecture, and how the different components of LaM-SLidE interact.
Figure 5: Expanded architectural overview. First Stage: The model is trained to reconstruct the encoded system, by querying the latent system representation by IDs. Second Stage: Latent flow-based model is trained to predict multiple masked future timesteps. The predicted system states are decoded by the frozen decoder.
@@ B.2 Identifier Assignment
We provide pseudocode for the identifier creation in Algorithm 1. This algorithm prevents the reuse of already assigned IDs, maintaining unique IDs across all entities. From a practical perspective we sample IDs randomly, so that all entity embeddings receive gradient updates.
Input : number of entities $N$; identifier pool size $|{\mathcal{I}}|$ where $N\leqslant|{\mathcal{I}}|$; embedding dimension $D_{u}$
@@
1 $\mathbf{U}\leftarrow$ empty matrix of size $N\times D_{u}$
@@ B.3 Encoder and Decoder
We provide pseudocode of the forward passes for encoding ($\mathcal{E}$) to and decoding ($\mathcal{D}$) from the latent system space of LaM-SLidE in Algorithm 2 and Algorithm 3 respectively. In general, encoder and decoder blocks follow the standard Transformer architecture [90] with feedforward and normalization layers. To simplify the explanation, we omitted additional implementation details here and refer readers to our provided source code.
@@ 4 $\mathbf{Q}\leftarrow\mathrm{Linear}(\mathbf{Z}_{\text{init}})$
Output : $[\mathbf{x},\mathbf{m}]\in\mathbb{R}^{D_{x}+D_{m}}$
@@ 5 $\mathbf{q}\leftarrow\mathrm{Linear}(\mathbf{u})$
For the decoding functionality presented in Algorithm 3, we made use of multiple specific decoder blocks depending on the actual task (e.g., for the molecules dataset, we use one decoder block for atom positions and one decoder block for atom types).
@@ B.4 Latent Flow Model
We provide pseudocode of the data prediction network $\bm{o}_{\theta}$ forward pass in Algorithm 4. The latent layer functionality is given by Algorithm 5. The architecture of the latent layers (i.e., our flow model) is based on Dehghani et al. [23], with the additional usage of adaptive layer norm (adaLN) [69] as also used for Diffusion Transformers [67]. The implementation is based on ParallelMLP block codes from Black Forest Labs [14], which was adopted to use it along the latent dimension as well as along the temporal dimension (see Figure 6). The velocity model is obtained via reparameterization as outlined in Section C.1.

### v4

@@ 3.4 Training Procedure
The training process follows a two-stage approach similar to latent diffusion models [77]. First Stage: We train the encoder $\mathcal{E}$ and decoder $\mathcal{D}$, to reconstruct entities from latent space using assigned IDs, see Figure˜2). Second Stage: We train the approximator $\mathcal{A}$ on the latent system representations produced by the frozen encoder $\mathcal{E}$ (details in Sections˜E.3 and 5).
@@ 4 Experiments
In order to evaluate LaM-SLidE we focus on three key aspects: (i) Robust generalization in diverse domains. We examine LaM-SLidE ’s generalization in different data domains in relation to other methods, for which we utilize tracking data from human motion behavior, trajectories of particle systems, and data from molecular dynamics (MD) simulations. (ii) Temporal adaptability. We evaluate temporal adaptability through various conditioning/prediction horizons, considering single/multi-frame conditioning and short/long-term predictions; (iii) Computational efficiency and scalability. Finally, we assess the inference time of LaM-SLidE , and evaluate performance with respect to model size. The subsequent sections show our key findings. For implementation details and additional results see App.˜E and G, for information on the datasets see Section˜E.1.
@@
For the MD experiments on peptides (Tetrapeptides), we use the Jensen-Shannon divergence (JSD) over the distribution of torsion angles, considering both backbone (BB) and side chain (SC) angles. To capture long temporal behavior, we use Time-lagged Independent Component Analysis (TICA) [70], focusing on the slowest components TIC 0 and TIC 1. To investigate metastable state transitions, we make use of Markov State Models (MSMs) [74, 65].
For inference time and scalability, we assess the number of function evaluations (NFE) and report the performance of our method for different model sizes. A comprehensive overview of the conditioning and prediction horizons for each experiment is shown in Table˜7.
@@ 4.1 Pedestrian Trajectory Forecasting (ETH-UCY)
For human motion behavior, we first consider the ETH-UCY dataset [68, 53], which provides pedestrian movement behavior, over five different scenes: ETH, Hotel, Univ, Zara1, and Zara2. We use the same setup as Han et al. [35], Xu et al. [94, 95], in which the methods obtain the first 8 frames as conditioning information and have to predict the next 12 frames. We report the minADE/minFDE, computed across 20 sampled trajectories.
@@ Results
As shown in Table˜2, our model performs competitively across all five scenes, achieving lower minFDE for Zara1 and Hotel scene, and the lowest minADE on the ETH scene. Notably, in contrast to the compared baselines, we did not create additional features like velocity and acceleration or imply any kind of connectivity between entities.
@@ 4.2 Basketball Player Trajectory Forecasting (NBA)
Table 3: Results on the NBA dataset: Compared methods have to predict player positions for 12 frames and are given the initial 8 frames as input. Results in terms of minADE/minFDE for the Rebounding and Scoring scenarios.
@@
Our second experiment examines human motion behavior in the context of basketball gameplay. We utilize the SportVU NBA movement dataset [99], which contains player movement data from the 2015-2016 NBA season. Each recorded frame includes ten player positions (5 for each team) and the ball position. We consider two different scenarios: a) Rebounding: containing scenes of missed shots; b) Scoring: containing scenes of a team scoring a basket. The interaction patterns – both in frequency and in their adversarial versus cooperative dynamics – exhibit different challenges as those in Section˜4.1. The evaluation procedure by Xu et al. [95], which we use, provides 8 frames as input conditioning and the consecutive 12 frames for prediction. The performance is reported by the minADE/minFDE metrics, which are computed across 20 sampled trajectories. For the reported metrics, only the trajectories of the players are considered.
@@ Results
As illustrated in Table˜3, our model shows robust performance across both scenarios, Rebounding and Scoring. For the Scoring scenario, LaM-SLidE achieves parity with SocialVAE [34] in terms of minADE and surpasses the performance in terms of minFDE. In the Rebounding scenario, we observe comparable but slightly lower performance of LaM-SLidE compared to SocialVAE. Trajectories sampled by our model are shown in Figure˜13.
@@ 4.3 N-Body System Dynamics (Particle Systems)
We evaluate our method across three distinct N-Body simulation scenarios: a) Charged Particles: comprising particles with randomly assigned charges +1/1+1/-1 interacting via Coulomb forces [51, 82]; b) Spring Dynamics: consisting of $N=5$ particles with randomized masses connected by springs with a probability $0.5$ between particle pairs [51]; and c) Gravitational Systems: containing $N=10$ particles with randomized masses and initial velocities governed by gravitational interactions [17]. For all three scenarios, we consider 10 conditioning frames and 20 prediction frames. In line with Han et al. [35], we use $3000$ trajectories for training and $2000$ trajectories for validation and testing, and we report ADE/FDE averaged over $K=5$ runs.
@@ Results
LaM-SLidE achieves the best performance in terms of ADE/FDE for the Charged Particles and Gravity scenarios and competitive second-rank performance in the Spring Dynamics scenario, see Table˜2. Unlike compared methods, LaM-SLidE achieves these results without computing intermediate physical quantities such as velocities or accelerations. We present sampled trajectories in Figure˜14.
@@ 4.4 Molecular Dynamics - Small Molecules (MD17)
In this experiment, we evaluate LaM-SLidE on the well-established MD17 [21] dataset, containing simulated molecular dynamics trajectories of 8 small molecules. The size of those molecules ranges from 9 atoms (Ethanol and Malonaldehyde) to 21 atoms (Aspirin). We use 10 frames as conditioning, 20 frames for prediction, and report ADE/FDE averaged over $K=5$ runs.
@@
Similar to Section˜4.3, we compared LaM-SLidE to: Tensor Field Network [89], Radial Field [52], SE(3)-Transformer [30], EGNN [82], EqMotion [94], GeoTDM [35], and SVAE [95].
@@ Results
The results in Table˜4 show the performance on the MD17 benchmark. LaM-SLidE achieves the lowest ADE/FDE of all methods and for all molecules. These results are particularly remarkable considering that: (1) our model operates without an explicit definition of molecular bond information, and (2) it surpasses the performance of all equivariant baselines, an inductive bias we intentionally omitted in LaM-SLidE .
@@
Table 4: Results on the MD17 dataset. Compared methods have to predict the atom positions of 20 frames, conditioned on 10 input frames. Results in terms of ADE/FDE, averaged over 5 runs.
@@ 4.5 Molecular Dynamics - Tetrapeptides (4AA)
To investigate LaM-SLidE on long prediction horizons, we utilize the Tetrapeptide dataset from Jing et al. [44], containing explicit-solvent molecular dynamics trajectories simulated using OpenMM [26]. The dataset comprises 3,109 training, 100 validation, and 100 test peptides. We use a single conditioning frame to predict 10,000 consecutive frames. The predictions are structured as a sequence of ten cascading 1,000-step rollouts, where each subsequent rollout is conditioned on the final frame of the previous. Note that, in contrast to the MD17 dataset, the methods predict trajectories of unseen molecules.
@@ Results
Table˜5 shows performance metrics of the methods (for details on those metrics see Section˜E.6). Figure˜16 shows the distribution of backbone torsion angles, and the free energy surfaces of the first two TICA components, for ground truth vs simulated trajectories. LaM-SLidE performs competitively with the current state-of-the-art method MDGen with respect to torsion angles, which is a notable achievement given that MDGen operates in torsion space only. With respect to the TICA and MSM metrics, LaM-SLidE even outperforms MDGen. Sampled trajectories are shown in Figure˜17.
Table 5: Results on the Tetrapeptide dataset: Columns denote the JSD between distributions of torsion angles (backbone ($B$), side-chain (SC), and all angles), the TICA, and the MSM metric.
@@ 4.6 Computational Efficiency and Scaling Behavior
To assess computational efficiency, we compare the NFEs of our model to the second-best method, GeoTDM. Our results show that LaM-SLidE requires up to 10x-100x fewer function evaluations depending on the domain. For a detailed discussion, see Section˜G.2.
We further assess the scalability of LaM-SLidE across different model sizes. Our results indicate that performance consistently improves with increasing parameter count, suggesting that our method benefits from larger model capacity and could potentially achieve even better results with additional computational resources. Comprehensive details of this analysis can be found in App.˜G.
4.7 Ablations
Identifier Pool Size
To understand how the size of the identifier pool affects the performance of the first stage model, we conducted additional experiments on the MD17 dataset. Maintaining a fixed number of update steps across different configurations results in a modest increase in the reconstruction error for larger pool sizes. We attribute this to the reduced number of updates per individual identifier embedding. A comprehensive analysis is provided in Section˜G.4.
Identifier Assignment
We assessed the sensitivity of our model to the specific entity-identifier assignment by evaluating five random assignments on the MD17 dataset. Results do not indicate a negative impact on the performance of our model. This demonstrates the robustness of our approach with respect to the specific assignment between entities and identifiers, see Section˜G.5 for more details.
Identifier Latent Utilization
We analyzed decoder attention patterns to understand how identifiers access latent information across different compression ratios. Our findings show that each identifier maintains a consistent addressing scheme across different molecule conformations, and the model adaptively utilizes attention heads based on the capacity of the latent space. In an overparameterized setting, the model favors single-head retrieval, while in the compressed setting utilizes both heads to avoid potential identifier collisions. See Section˜G.6 for detailed analysis.
@@ 5 Discussion
Our experiments indicate that our architecture applies to a diverse set of problems; however, a few limitations provide opportunities for future improvement. While our current approach successfully allows for compressing entities with beneficial reconstruction performance, our experiments indicate a tradeoff between the number of latent space vectors to encode system states and the performance of our latent model; for more details, see Section˜G.3.
@@ Conclusion
We have introduced LaM-SLidE, a novel approach for modeling spatial dynamical systems that consist of a variable number of entities within a fixed-size latent system representation, where assignable identifiers enable the traceability of individual entities. Across diverse domains, LaM-SLidE matches or exceeds specialized methods and offers promising scalability properties. Its minimal reliance on prior knowledge makes it suitable for many tasks, suggesting its potential as a foundational architecture for dynamical systems.
@@ Acknowledgments
The ELLIS Unit Linz, the LIT AI Lab, and the Institute for Machine Learning are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We thank NXAI GmbH, Audi AG, Silicon Austria Labs (SAL), Merck Healthcare KGaA, GLS (Univ. Waterloo), TÜV Holding GmbH, Software Competence Center Hagenberg GmbH, dSPACE GmbH, TRUMPF SE + Co. KG. We further thank Niklas Schmidinger, Anna Zimmel, Benedikt Alkin, and Arturs Berzins for helpful discussions and feedback.
@@ Appendix A Notation
continuous time $\hat{t} \mathbb{R}$
overall number of (sampled) time steps $T$ $\mathbb{N}$
number of observed time steps (when predicting later ones) $T_{o}$ $\mathbb{N}$
number of future time steps (prediction horizon) $T_{f}=T-T_{o}$ $\mathbb{N}$
time index for sequences of time steps $t$ $\mathbb{N}$
@@
spatial entity dimensionality $D_{x}$ $\mathbb{N}$
entity feature dimensionality $D_{m}$ $\mathbb{N}$
@@
identifier representation dimensionality $D_{u}$ $\mathbb{N}$
number of latent vectors $L$ $\mathbb{N}$
latent vector dimensionality $D_{z}$ $\mathbb{N}$
@@ B.1 Architecture Overview in Detail
Figure˜5 shows an expanded view of our architecture, and how the different components of LaM-SLidE interact.
Figure 5: Expanded architectural overview. First Stage: The model is trained to reconstruct the encoded system by querying the latent system representation by IDs. Second Stage: Latent flow-based model is trained to predict multiple masked future timesteps. The predicted system states are decoded by the frozen decoder.
@@ B.2 Identifier Assignment
We provide pseudocode for the identifier creation in Algorithm˜1. This algorithm prevents the reuse of already assigned IDs, maintaining unique IDs across all entities. From a practical perspective, we sample IDs randomly so that all entity embeddings receive gradient updates.
Input : number of entities NN; identifier pool size $|{\mathcal{I}}|$ where $N\leqslant|{\mathcal{I}}|$; embedding dimension $D_{u}$
@@
1
[…]
@@ B.3 Encoder and Decoder
We provide pseudocode of the forward passes for encoding ($\mathcal{E}$) to and decoding ($\mathcal{D}$) from the latent system space of LaM-SLidE in Algorithm˜2 and Algorithm˜3 respectively. In general, encoder and decoder blocks follow the standard Transformer architecture [91] with feedforward and normalization layers. To simplify the explanation, we omitted additional implementation details here and refer readers to our provided source code.
@@ 4 $\mathbf{Q}\leftarrow\mathrm{Linear}(\mathbf{Z}_{\text{init}})$
Output : $\mathbf{x},\mathbf{m}]\in\mathbb{R}^{D_{x}+D_{m}}$
@@ 5 $\mathbf{q}\leftarrow\mathrm{Linear}(\mathbf{u})$
For the decoding functionality presented in Algorithm˜3, we made use of multiple specific decoder blocks depending on the actual task (e.g., for the molecules dataset, we use one decoder block for atom positions and one decoder block for atom types).
@@ B.4 Latent Flow Model
We provide pseudocode of the data prediction network $\bm{o}_{\theta}$ forward pass in Algorithm˜4. The latent layer functionality is given by Algorithm˜5. The architecture of the latent layers (i.e., our flow model) is based on Dehghani et al. [23], with the additional usage of adaptive layer norm (adaLN) [69] as also used for Diffusion Transformers [67]. The implementation is based on ParallelMLP block code from Black Forest Labs [14], which was adapted to use it along the latent dimension as well as along the temporal dimension (see Figure˜6). The velocity model is obtained via reparameterization as outlined in Section˜C.1.